# A First Look at GPT Apps: Landscape and Vulnerability The original version was uploaded on Feb 23, 2024

### v2

@@ Abstract.
Following OpenAI’s introduction of GPTs, a surge in GPT apps has led to the launch of dedicated LLM app stores. Nevertheless, given its debut, there is a lack of sufficient understanding of this new ecosystem. To fill this gap, this paper presents a first comprehensive longitudinal (5-month) study of the evolution, landscape, and vulnerability of the emerging LLM app ecosystem, focusing on two GPT app stores: GPTStore.AI and the official OpenAI GPT Store. Specifically, we develop two automated tools and a TriLevel configuration extraction strategy to efficiently gather metadata (i.e., names, creators, descriptions, etc.) and user feedback for all GPT apps across these two stores, as well as configurations (i.e., system prompts, knowledge files, and APIs) for the top 10,000 popular apps. Our extensive analysis reveals: (1) the user enthusiasm for GPT apps consistently rises, whereas creator interest plateaus within three months of GPTs’ launch; (2) nearly 90% system prompts can be easily accessed due to widespread failure to secure GPT app configurations, leading to considerable plagiarism and duplication among apps. Our findings highlight the necessity of enhancing the LLM app ecosystem by the app stores, creators, and users.
@@ 1. Introduction
Large Language Models (LLMs) such as OpenAI’s GPT-4 (Achiam et al., 2023), Google’s Gemini (Masalkhi et al., 2024), and Meta’s LLaMa (Touvron et al., 2023), which are pre-trained on vast datasets, have demonstrated remarkable language understanding and reasoning capabilities. The advent of LLMs has catalyzed the development of powerful applications by leveraging simple textual prompts and external materials, further influencing various domains, including finance, healthcare, education, and entertainment. As of Jan. 2024 (OpenAI, [n. d.]), the proliferation of GPT-based applications has surpassed three million.  [^1] These authors contributed equally to this paper.
@@
Investigating app stores is crucial for understanding the development of apps in real-world scenarios, as it provides insights into user engagement, technological and market trends, app dynamics, and privacy concerns. For instance, a large amount of research work has been conducted on the app ecosystems of Android (Viennot et al., 2014; Wang et al., 2019; Potharaju et al., 2017) and iOS app stores (Lin et al., 2021; Erić et al., 2014), examining the characteristics and evolution of mobile apps (Seneviratne et al., 2015; Petsas et al., 2017; Wang et al., 2019; Carbunar and Potharaju, 2015), user behavior and preferences (Fu et al., 2013; Khalid et al., 2014; Li et al., 2016), mobile app recommendations (Karatzoglou et al., 2012; Yin et al., 2013; Liu et al., 2016), and the comparative analysis of different app stores (Liu et al., 2015; Trecca et al., 2021; Ali et al., 2017). Conducting studies on apps and app stores could lead to higher user satisfaction, enhanced app visibility and user trust (McIlroy et al., 2016; Sällberg et al., 2023).
However, the LLM app ecosystem remains unexplored, with no comprehensive analysis conducted to date. Compared to mobile apps, LLM apps are easier to build – everyone can create an LLM app without solid background knowledge or significant server and storage resources. Especially, the great simplicity of creating LLM apps comes with a frangibility – LLMs are known to be vulnerable to attacks (Tao et al., 2023; Liu et al., 2023), increasing the risk of data leaks. Therefore, a thorough analysis of the characteristics and vulnerabilities of LLM apps is crucial. Such insights are essential for providing valuable guidance to the app market and developers, assisting in the development and protection of user-centric GPT apps that positively impact society.
To fill this gap, our study conducts a longitude measurement of GPT app stores, which to our knowledge, is the first large-scale analysis of the LLM app ecosystem. Our research encompasses two primary objectives: 1) mapping out the landscape of GPT apps within these stores and 2) investigating potential vulnerabilities of GPT apps and plagiarism in GPT app stores. For the first goal, we take a 5-month monitor of two app stores, GPTStore.AI and OpenAI GPT Store, capturing metadata and user feedback. We conduct weekly configuration extraction of GPT apps to analyze their characteristics and dynamics, including system prompts, knowledge files, and APIs. Our study focuses on answering the following research questions:
@@
Our research faces several key challenges in answering the above research questions. First, there is no public dataset with comprehensive GPT app information. Thus, it is essential to build an efficient web scraping tool for data capturing and longitudinal monitoring. Second, unlike studies of mobile app stores that focus primarily on static information (rankings, description, etc.), understanding GPT apps requires interactions with them (e.g., chatbot). Consequently, building an automated tool for large-scale interaction with GPT apps is crucial. Third, to examine apps’ vulnerability, it is difficult to develop effective configuration extraction methods that can extract comprehensive information from these apps.
Landscape: To analyze the landscape of GPT app stores, we develop a web scraping tool that automatically captures webpages and extracts metadata and user feedback from all GPT apps available in the two app stores we investigate.
@@
We collect a five-month dataset consisting of three types of data: app metadata, user feedback, and GPT configurations (detailed in §3.2). The metadata and user feedback come from the longitudinal monitoring of web information, while GPT configurations are obtained through the configuration extraction of the top 10,000 GPT apps, as well as weekly capture of the top 500 GPT apps. Through analyzing these data, we uncover some valuable insights into the dynamics and evolution within the GPT app ecosystem. We highlight our key findings below:
$\bullet$ Enthusiasm for creating and interacting with GPT apps initially peaks but has recently become less appealing to creators. We observe significant enthusiasm among users and creators for interacting with and creating GPT apps, as evidenced by the rapid growth to more than 97,049 GPT apps and 30,657 creators documented in GPTStore.AI within just five months. However, this initial surge in the number of apps and creators levels off after the first three months, with only about 200 apps being updated or added each week thereafter. This slowdown indicates a decline in creator enthusiasm for developing new GPT apps, possibly due to challenges in maintaining innovation or a lack of sufficient incentives.
$\bullet$ GPT apps usage exhibits clear power law, where a small number of apps account for most user interactions. The GPT app store displays a pattern similar to other mobile app stores  (Carbunar and Potharaju, 2015), adhering to the Pareto Principle (Sanders, 1987)—where approximately 80% of effects come from 20% of causes. However, in the GPT app store, this imbalance is even more marked. The top 1% of apps are responsible for 90% of daily conversations. Specifically, by the end of April, the top five apps account for 25% of daily overall conversations, while the top 50 apps constitute 60%. Similarly, only 18% of apps have user feedback. This shows a significant concentration of user engagement within a small subset of available apps, indicating that most apps are underutilized.
$\bullet$ The high success rate of configuration extraction of GPT apps exposes their vulnerabilities and indicates a lack of sufficient protective mechanisms. Through configuration extraction of the top 10,000 GPT apps, we successfully accessed system prompts, the names of knowledge files, and the contents of knowledge files from 90%, 88%, and 12.7% of them, respectively. Even with recent protective measures possibly implemented by OpenAI, our success rate for acquiring system prompts remains above 77%. This high success rate highlights the vulnerability of GPT apps, potentially exposing creators’ private information and intellectual property.
$\bullet$ Notable instances of plagiarism within the GPT app store have raised copyright issues. The ease of accessing GPT configurations has led to significant duplication and plagiarism within the GPT app store. In particular, we identify over 705 GPT apps among the top 10,000 that share nearly identical names and descriptions, which may confuse users when searching with specific keywords. We also notice that there are many cases where multiple GPT apps share identical or very similar system prompts, yet differ in names, descriptions, and creators. These GPT apps often appear at the top of search results and are frequently chosen by users. This not only raises copyright issues but also limits the diversity of options available within the GPT app ecosystem.
Our findings have significant implications for GPT users, creators, and app stores. (1) For GPT app stores: First, unlike traditional closed-source mobile apps uploaded to mobile app stores, the system prompts and knowledge files used to construct GPT apps are transparent to the GPT store. Given the vulnerability of GPT apps, we recommend that app stores implement continuous monitoring to notify GPT creators of potential leaks and duplication issues regarding GPT configurations. Second, to ensure the sustainable development of GPT apps, the GPT store should promptly introduce financial incentives to attract creators to this market. (2) For GPT creators: They should be aware of the importance of protecting GPT configurations. Currently, it is better not to include proprietary information in system prompts and knowledge files, as protective instructions cannot guarantee complete security. A possible solution is to move sensitive data to external API calls, although this increases the developers’ burden. (3) For users: We highly advise them to actively provide feedback and ratings. Such engagement is crucial for improving the quality and reliability of apps within the ecosystem.
@@
Our contributions are summarized as follows: (1) We are the first to conduct extensive monitoring and analysis of the LLM app ecosystem. (2) We introduce a novel TriLevel GPT app configuration extraction approach. (3) Our research provides a detailed dataset on GPT apps, which we intend to open-source if our paper is accepted.
@@
Our study does not raise any ethical issues. The details can be found in Appendix A.
@@ 2.1. LLM App Stores
LLM Apps: LLM apps are software programs that utilize the capabilities of LLMs along with a variety of external skills to perform diverse tasks. Common examples include chatbots, text-based games, image generators, etc. The emergence of LLM app stores addresses the increasing need for a centralized platform where users can discover, explore, and interact with a range of LLM apps. In LLM app stores, apps are presented with thorough descriptions and categorizations, which simplifies the discovery process and enhances user engagement.
LLM App Stores: Multiple LLM app stores exist, each playing a significant role in demonstrating the potential of LLMs and facilitating their adoption for various applications and scenarios. For instance, OpenAI’s official GPT store exemplifies such a platform, offering a variety of GPT apps built on its GPT model. Poe (Poe, [n. d.]), powered by Quora, is a platform that allows users to discover, interact, and build various bots from third-party LLMs and developers, such as GPT from OpenAI, Claude from Anthropic, and Gemini from Google. Another notable LLM app store is FlowGPT (FlowGPT, [n. d.]) which builds a vibrant community that supports both developers and users, further exemplifying the diversity and utility of LLM app stores.
@@ 2.2. GPT App Stores
Our study focuses on one of the largest LLM apps – GPT apps, and corresponding GPT app stores. We aim to track their development and evolution over a period of five months, identifying potential vulnerabilities within these stores and GPT apps themselves. This research offers crucial insights into the LLM app ecosystem and underscores the importance of centralized platforms in fostering the adoption and advancement of LLM apps.
@@
GPT Apps Development: OpenAI enables the customization of GPT applications by integrating instructions, additional knowledge, and a blend of inherent and external capabilities. As depicted in Figure 1, the architecture of a GPT application consists of three configurations.
$\bullet$ Instructions: This section includes essential details such as the application’s name, description, logo, system prompts, prompt starter, etc.
$\bullet$ Knowledge: Users can upload external knowledge files to produce contextually accurate responses, ensure the fairness of the model’s outputs, and tailor the user experience.
$\bullet$ Actions: This includes built-in features like image generation through DALL·E, web browsing, and code interpretation using Python, along with the ability to integrate third-party services through external APIs.
Once the app configuration is complete, we can upload it and make it publicly available. There is no review process required before release.
GPT App Stores: OpenAI offers an official GPT store to enhance user navigation and engagement. This store organizes GPT apps into seven distinct categories: Dall·E, Writing, Productivity, Research & Analysis, Programming, Education, and Lifestyle, and it features the top 12 apps in each category. However, it does not offer a comprehensive dataset of GPT apps. To fill this gap, we examine GPTStore.AI, a repository of GPT apps, which categorizes apps into 50 distinct categories and provides access to all available apps. Additionally, GPTStore.AI offers detailed metadata to facilitate further exploration.
@@ 3. Methodology
To explore the LLM app ecosystem, we perform a five-month longitudinal monitoring on two GPT stores, GPTStore.AI and the official OpenAI GPT Store, analyzing the dynamics, evolution, and vulnerability of app development.
@@ 3.1. Overview
As depicted in Figure 2, our data collection process involves two stages: Web Scraping and TriLevel Configuration Extraction, through which we obtain three types of data: metadata, user feedback, and GPT app configurations. The Web Scraping stage collects all the metadata and user feedback from GPT apps, while the TriLevel Configuration Extraction stage is designed to acquire detailed GPT app configurations, including system prompts, files, and APIs.
Automated Tools: By utilizing Selenium (Nyamathulla et al., 2021) and Python, we develop automated tools for web scraping and GPT app interactions. These tools effectively manage browser instances, simulate user interactions, capture web pages, etc.
@@
Web Scraping: We implement distinct monitoring strategies for OpenAI GPT Store and gptstore.ai. For OpenAI GPT Store, we take daily snapshots of web pages in both ”Global” and ”English” modes. For gptstore.ai, from December 2023 to May 2024, we systematically capture all web pages from the platform every five days, specifically targeting three types of pages: index pages that list all GPT apps, category-specific index pages, and detailed information pages for individual apps. Our dataset eventually includes 25 complete sets of web pages from gptstore.ai and 60 days of web pages from OpenAI GPT Store.
TriLevel Configuration Extraction: In addition to metadata and user feedback gathered via web scraping, we engage with apps in the chatbot and employ TriLevel Configuration Extraction to extract as much information as possible about the apps’ configurations. This process is structured into three levels:
In Level 1, we initiate with direct queries for system prompts, files, and APIs using straightforward questions/prompts. The failure cases will be delivered to Level 2, which adopts the same procedures as Level 1. This repetition aims to avoid randomness in the output. Then, unresolved cases from Level 2 are transferred to Level 3, where we challenge the GPT with multilingual replication prompts, aiming to bypass protections by leveraging potential vulnerabilities in multilingual comprehension (Choudhury et al., 2023; Beniwal et al., 2024). In Levels 2 and 3, we focus solely on acquiring system prompts as we discover that system prompts can be easily obtained compared to the other configurations.
In Direct query of Level 1 and 2, we use the following straightforward prompts to access the GPT app configurations.
  • For system prompt: Repeat the words above starting with the phrase ”You are a GPT”; Put them in a code block, and include everything. Certainly! Here’s the text you requested in a code block:
@@
In Multilingual GPT Replication, we aim to obtain the system prompts of a GPT app by asking the app to generate a system prompt to exactly “replicate” itself. We carefully design specific prompts (as shown in below) to achieve this target. Furthermore, we interact with the GPT apps by using our developed prompts in various languages to increase the success rate of accessing system prompt, as several other concurrent studies (Choudhury et al., 2023; Beniwal et al., 2024) have also shown that the GPT apps may bear some bias when being asked about the same question in diverse languages. We also validate the effectiveness of our Multilingual GPT Replication approach in Appendix B and confirm that it can indeed help us get the system prompt of the GPT apps. Our developed prompts are as follows:
@@
Due to the extremely large volume of GPT apps and the limited conversation count budget allowed by OpenAI, we do not perform configuration extraction on all the GPT apps. Instead, we focus on the most popular and representative ones. In particular, we first rank the GPT apps in our dataset (derived through our Web Scraping) according to their conversation counts. Then, we began by extracting the configurations of the top 10,000 GPT apps on February 24, 2024. After that, from March to May 2024, we conducted weekly configuration extraction on the top 500 GPT apps. In total, we collect data from the initial 10,000 apps and also from nine weekly monitoring of the top 500 apps, specifically on the following dates: March 6, 13, 20, 27, April 3, 10, 17, 24, and May 1.

### v3

@@ Abstract.
Following OpenAI’s introduction of GPTs, a surge in GPT apps has led to the launch of dedicated LLM app stores. Nevertheless, given its debut, there is a lack of sufficient understanding of this new ecosystem. To fill this gap, this paper presents a first comprehensive longitudinal (10-month) study of the evolution, landscape, and vulnerability of the emerging LLM app ecosystem, focusing on three GPT app stores: GPTStore.AI and GPTsHunter and the official OpenAI GPT Store. Specifically, we develop two automated tools and a TriLevel configuration extraction strategy to efficiently gather metadata (names, creators, descriptions, etc.) and user feedback for all GPT apps across these three stores, as well as configurations (system prompts, knowledge files, and APIs) for the top 10,000 popular apps. Our extensive analysis reveals: (1) The user enthusiasm for GPT apps consistently rises, whereas creator interest plateaus within five months of GPTs’ launch. (2) Nearly 90% of system prompts can be easily accessed due to widespread failure to secure GPT app configurations, leading to considerable plagiarism and duplication among apps. Our findings highlight the necessity of enhancing the LLM app ecosystem by the app stores, creators, and users.
copyright: nonejournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Large Language Models (LLMs) such as OpenAI’s GPT-4 (Achiam et al., 2023), Google’s Gemini (Masalkhi et al., 2024), and Meta’s LLaMa (Touvron et al., 2023), which are pre-trained on vast datasets, have demonstrated remarkable language understanding and reasoning capabilities. The advent of LLMs has catalyzed the development of powerful applications by leveraging simple textual prompts and external materials, further influencing various domains, including finance, healthcare, education, and entertainment. As of Jan. 2024 (OpenAI, [n. d.]), the proliferation of GPT-based applications has surpassed three million.
@@
The study of LLM app stores is crucial for understanding the real-world development of web LLM applications, as it offers insights into user engagement, technological trends, market dynamics, and privacy concerns. Similar research has been extensively conducted on Android (Viennot et al., 2014; Wang et al., 2019; Potharaju et al., 2017) and iOS app ecosystems (Lin et al., 2021; Erić et al., 2014), where researchers have explored app characteristics and evolution (Seneviratne et al., 2015; Petsas et al., 2017; Wang et al., 2019; Carbunar and Potharaju, 2015), user behavior and preferences (Fu et al., 2013; Khalid et al., 2014; Li et al., 2016), app recommendation systems (Karatzoglou et al., 2012; Yin et al., 2013; Liu et al., 2016), and comparative analysis of app stores (Liu et al., 2015; Trecca et al., 2021; Ali et al., 2017). Such investigations can lead to higher user satisfaction, improved app visibility, and increased user trust (McIlroy et al., 2016; Sällberg et al., 2023).
Relevance to Web: LLM apps and their ecosystems are fundamentally web-based, relying on web infrastructure for core operations. Without web connectivity, LLM apps cannot function, as they depend on real-time data processing and retrieval from remote servers. Additionally, LLM app stores, where users discover, download, and interact with these apps, operate primarily as web platforms, providing access to a wide range of LLM services. These platforms are essential gateways for accessing and engaging with LLM apps and services.
However, the LLM app ecosystem remains largely unexplored, with no comprehensive longitudinal analysis conducted to date. Some studies focus on limited app metadata (Zhao et al., 2024b, a) or use outdated, small-scale datasets (Su et al., 2024), while others provide analyses without real-world validation (Zhao et al., 2024a). Beside, unlike mobile apps, LLM-based applications are much easier to develop—anyone can create an LLM app online without extensive technical knowledge or significant network and storage resources. However, this ease of development comes with increased vulnerability, as LLMs are known to be susceptible to attacks (Tao et al., 2023; Liu et al., 2023), raising concerns about data leaks. Therefore, a thorough analysis of the characteristics and vulnerabilities of LLM apps is essential to provide insights into their development and deployment. This would offer valuable guidance to app markets and creators, while promoting the creation of secure, user-centric GPT apps that positively impact society.
To fill this gap, our study conducts a longitude measurement of GPT app stores, which, to our knowledge, is the first large-scale analysis of the LLM app ecosystem. This research focuses on two primary objectives: 1) mapping out the landscape of GPT apps within these web stores, and 2) investigating potential vulnerabilities and plagiarism in GPT app stores. For the first goal, we perform a 10-month monitoring of three app stores—GPTStore.AI, GPTshunter and official OpenAI GPT Store—capturing metadata and user feedback. We conduct weekly configuration extraction of GPT apps to analyze their characteristics and dynamics, including system prompts, knowledge files, and APIs. Our study aims to address the following research questions:
@@
Our research faces several key challenges in answering the above research questions. First, there is no public dataset that provides comprehensive and long-term GPT app information. Thus, it is essential to build an efficient web scraping tool for data capturing and longitudinal monitoring. Second, unlike studies of mobile app stores that focus primarily on static information (rankings, description, etc.), understanding GPT apps requires interactions with them (e.g., chatbot). Consequently, building an automated tool for large-scale interaction with GPT apps is crucial. Third, to examine apps’ vulnerability, it is difficult to develop effective configuration extraction methods that can extract comprehensive information from these apps.
Landscape: To analyze the landscape of GPT app stores, we develop a web scraping tool that automatically captures webpages and extracts metadata and user feedback from all GPT apps available in the three app stores we investigate.
@@
We collect a ten-month dataset consisting of three types of data: app metadata, user feedback, and GPT configurations (detailed in §3.2). The metadata and user feedback come from the longitudinal monitoring of web information, while GPT configurations are obtained through the configuration extraction of the top 10,000 GPT apps, as well as weekly capture of the top 500 GPT apps. By analyzing these data, we uncover valuable insights into the dynamics and evolution of the GPT app ecosystem.
Our study shows that after an initial surge in GPT app creation, growth significantly slows after five months, indicating declining creator enthusiasm. User engagement follows a power-law distribution, with the top 1% of apps driving 90% of daily interactions, leaving most apps underutilized. We also identify vulnerabilities in GPT app configurations, with system prompts, knowledge file names, and file contents successfully extracted from 90%, 88%, and 12.7% of apps, respectively, posing risks to creators’ data. Additionally, over 705 apps have nearly identical names and descriptions, raising concerns about plagiarism and copyright and limiting diversity in the app ecosystem.
Thus, we recommend that GPT app stores implement continuous monitoring to alert creators of potential leaks or duplication, along with offering financial incentives to drive growth. Creators should avoid placing proprietary information in system prompts and consider using external APIs to safeguard sensitive data. Financial incentives could also motivate creators to enhance their apps. Additionally, users are encouraged to provide active feedback to help improve app quality and security.
@@
Our contributions are as follows: (1) We conduct the first long-term, large-scale monitoring and analysis of the LLM app ecosystem. (2) We introduce a novel TriLevel GPT app configuration extraction approach. (3) Our research provides a comprehensive dataset on GPT apps, which we plan to open-source upon paper acceptance.
@@
Our study does not encounter ethical issues. The detailed discussion can be found in Appendix A.
@@ 2.1. LLM App Stores
LLM apps use large language models (LLMs) and external tools for tasks like chatbots, text-based games, and image generation. LLM app stores provide a centralized platform where users can discover, explore, and interact with these apps, with detailed descriptions and categories to make navigation easier and boost engagement. Notable examples include OpenAI’s GPT store, where users can interact with and rate apps; Poe (Poe, [n. d.]) by Quora, which supports bots from third-party LLMs; FlowGPT (FlowGPT, [n. d.]), which builds a community for developers and users, highlighting the variety and functionality of LLM apps.
@@ 2.2. GPT App Stores
Our study focuses on GPT apps and their stores, aiming to provide key insights into the LLM app ecosystem and highlight the importance of centralized platforms in promoting LLM app growth. Several GPT app stores exist, such as GPTStore.AI (GPTStore.ai, [n. d.]), GPTshunter (Hunter, [n. d.]), GIPETIES (Gipeties, [n. d.]), GPTsdex (GPTsdex, [n. d.]), and Epic GPT Store (Store, [n. d.]). In January 2024, OpenAI launched its official GPT store. These platforms improve user navigation and engagement through direct access links and detailed app descriptions.
@@
GPT Apps Development: OpenAI provides a framework for customizing GPT apps by integrating instructions, additional knowledge, and built-in or external functionalities. As depicted in Figure 1, GPT app configuration consists of three main components:
$\bullet$ Instructions: Includes the app’s name, description, logo, system prompts, etc.
$\bullet$ Knowledge: Users can upload external knowledge files to improve response accuracy, fairness, and the user experience.
$\bullet$ Actions: Offers features like DALL·E for image generation, web browsing, Python code interpretation, and external APIa.
Once configured, apps can be uploaded and made publicly available. No review process is required before release.
@@ 3. Methodology
To explore the LLM app ecosystem, we perform a ten-month longitudinal monitoring on three GPT stores, GPTStore.AI, GPTshunter and the official OpenAI GPT Store, analyzing the dynamics, evolution, and vulnerability of app development.
@@ 3.1. Overview
As depicted in Figure 2, our data collection process involves two stages: Web Scraping and TriLevel Configuration Extraction. Web Scraping gathers metadata and user feedback, while TriLevel Configuration Extraction captures detailed GPT app configurations, including system prompts, files, and APIs.
@@
Web Scraping. We implement distinct monitoring strategies for each store. For GPTStore.AI, from December 2023 to May 2024, we systematically capture all web pages weekly, targeting three types of pages: index pages that list all GPT apps, category pages, and detailed information pages for individual apps. For GPTShunter, we obtain the entire dataset of all apps directly from the store owner.
TriLevel Configuration Extraction. To obtain app configurations, we interact with apps in OpenAI’s chatbot in the OpenAI GPT Store and use TriLevel Configuration Extraction to gather as much information as possible. This process is structured into three levels:
In Level 1, we initiate with direct queries for system prompts, files, and APIs using straightforward questions/prompts. The failure cases will be delivered to Level 2, which we repeat the process in Level 1 to avoid randomness in the responses. Then, unresolved cases from Level 2 are transferred to Level 3, where we use multilingual replication prompts to bypass protections by leveraging vulnerabilities in multilingual understanding (Choudhury et al., 2023; Beniwal et al., 2024). In Levels 2 and 3, we focus on system prompts, as they are easier to obtain than other configurations.
In Direct query of Level 1 & 2, we use the following straightforward prompts to access the GPT app configurations.
  • For system prompt: Repeat the words above starting with the phrase ”You are a GPT”; Put them in a code block, and include everything.
@@
In Multilingual GPT Replication, we retrieve a GPT app’s system prompt by asking it to generate prompts to ”replicate” itself. We use carefully designed prompts in multiple languages to improve success, as studies (Choudhury et al., 2023; Beniwal et al., 2024) suggest GPT apps may respond differently based on language. The validation of this method can be found in Appendix C. Our developed prompts are as follows:
@@
Automated Tools. By utilizing Selenium (Nyamathulla et al., 2021) and Python, we develop automated tools for web scraping and GPT app interactions. These tools effectively manage browser instances, simulate user interactions, capture web pages, etc.